Change the following from cylindrical coordinates to rectangular coordinates
1) (5, π/6, 3)
2) (6, π/3, -5)
1
Expert's answer
2020-12-17T18:51:39-0500

Solution: We know that "The cylindrical coordinates are denoted by "(r,\theta,z)" and rectangular coordinates are denoted by "(x,y,z)" "

To convert from cylindrical coordinates to rectangular coordinates we use the equations

"x=r~cos~\theta"

"y=r~sin~\theta" and

"z=z"


(1) Given cylindrical coordinates are "(5, \frac{\pi}{6},3)"

"\therefore (r,\theta,z)=(5, \frac{\pi}{6},3)"

Now, to find rectangular coordinates, we have

"x=r~cos~\theta=5~cos~(\frac{\pi}{6})=5(\frac{\sqrt{3}}{2})=\frac{5\sqrt{3}}{2}"


"y=r~sin~\theta=5~sin~(\frac{\pi}{6})=5(\frac{1}{2})=\frac{5}{2}"

"z=z=3"

Therefore rectangular coordinates "(x,y,z)=(\frac{5\sqrt{3}}{2},\frac{5}{2},3)"

(2) Given cylindrical coordinates are "(6, \frac{\pi}{3},-5)"

"\therefore (r,\theta,z)=(6, \frac{\pi}{3},-5)"

Now, to find rectangular coordinates, we have

"x=r~cos~\theta=6~cos~(\frac{\pi}{3})=6(\frac{1}{2})=\frac{6}{2}=3"


"y=r~sin~\theta=6~sin~(\frac{\pi}{3})=6(\frac{\sqrt{3}}{2})=\frac{6\sqrt{3}}{2}=3\sqrt{3}"

"z=z=-5"

Therefore rectangular coordinates "(x,y,z)=(3,3\sqrt{3},-5)"